if the area of a triangle is 54 square inches and the base is 12 inches, then what is the height?
Answer by nerdybill(5411)

if the area of a triangle is 54 square inches and the base is 12 inches, then what is the height?
.
Remember, the area of any triangle is:
(1/2)bh
where
b is length of the base
h is length of the height
.
the problem gives us
area as 54 sq inches
base as 12 inches
.
plugging it in we get:
(1/2)bh = area
(1/2)(12)h = 54
6h = 54
h = 9 inches